ON THE ADVANTAGES OF MATRIX-ISOLATION VIBRATIONAL CIRCULAR DICHROISM SPEC-TROSCOPY

\underline{GYÖRGY TARCZAY}, GÁBOR MAGYARFALVI, and ELEMÉR VASS, Institute of Chemistry, Eötvös University, H-1518 Budapest 112, P.O. Box 32, Hungary.

It is shown on the examples of R-2-amino-1-propanol ^{a} and N-acetyl-L-alanyl-N-methylamide ^{b} that matrix isolation vibrational circular dichroism spectroscopy is a useful method for simultaneous determination of the conformational distribution and absolute configuration of molecules with multiple conformers capable of strong intermolecular interactions, especially hydrogen-bonds. The technique could also be utilized for the characterization of reactive chiral species or weakly bound complexes of chiral molecules.

aG. Tarczay, G. Magyarfalvi and E. Vass Angew. Chem., in press 2006.

bG. Pohl, G. Magyarfalvi, E. Vass, A. Perczel, G. Tarczay to be submitted.